John was studying how far the UMN students travelled to take the class orn campus. Let W be the distance travelled (measured in miles). A sample of size 25 was collected. He calculated the following statistics: W -3.8 Calculate the sample variance. Please round the answer to the nearest hundredth.
Answer #1

from above sum of square SS =23

and sample size n=25

here as we know that sample variance =SS/(n-1) =23/(25-1)=0.96